OpenStudy (anonymous):

Simplify the 6th root of 128b^2

OpenStudy (anonymous):

2 6th root 2b^2

OpenStudy (anonymous):

Is it 128*(b^2) ?

OpenStudy (anonymous):

\[\sqrt[6]{128 b^2}=2 \left(\sqrt[6]{2} \sqrt[3]{b}\right) \]

OpenStudy (anonymous):

How did you come up with that?

OpenStudy (triciaal):